OpenStudy (anonymous):

Algebra 2 help please

OpenStudy (anonymous):

Felix exclaims that his quadratic with a discriminant of −1 has no real solutions. Felix then puts down his pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Felix, in calm and complete sentences, how to find the solutions, even though they are not real.

OpenStudy (kirbykirby):

One simple equation would be \(x^2+1=0\)

OpenStudy (anonymous):

wouldn't have to be \[x ^{2} + bx + c\]?

OpenStudy (anonymous):

@kirbykirby

OpenStudy (kirbykirby):

as long as you have an \(x^2\) term (and that is the highest power in your polynomial), then it is a quadratic function. You have have here that \(a=1, b=0, c=1\)

OpenStudy (anonymous):

Oh alright, so would I have to find the discriminant?

OpenStudy (kirbykirby):

Notice that \(a\) can be any number except 0, because then \(x^2\) disappears

OpenStudy (anonymous):

Yeah I understand that now

OpenStudy (kirbykirby):

Well you can proceed as usual using the quadratic formula. Do you know about \(i^2=-1\)? If not you can just keep your solution with a negative square root

OpenStudy (anonymous):

I do know what i^2 is. So if I had x^2+2x+3 the discriminant would be -8?

OpenStudy (kirbykirby):

Yes indeed

OpenStudy (anonymous):

How would I find the solutions of those?

OpenStudy (kirbykirby):

Let's consider the simpler one first perhaps: \(x^2+1=0\) Then by the quadratic formula, \[ s=\frac{-0\pm\sqrt{0-4(1)(1)}}{2(1)}=\pm\frac{\sqrt{-4}}{2}\] Sine \(\sqrt{-1}=i\), then you get above that \[x=\pm\frac{2i}{2}=\pm i \]

OpenStudy (kirbykirby):

Since*

OpenStudy (kirbykirby):

And sorry I don't know why I wrote \(s=...\), I meant to write x

OpenStudy (anonymous):

But isn't the discriminant formula b^2 -4ac?

OpenStudy (kirbykirby):

yes, b=0, a=1, c=1

OpenStudy (kirbykirby):

0^2=0

OpenStudy (anonymous):

That's the quadratic formula..

OpenStudy (kirbykirby):

the discriminant is found within the quadratic formula

OpenStudy (kirbykirby):

Discriminant: \[ \Delta=b^2-4ac\] Quadratic formula: \[ \frac{-b\pm \sqrt{b^2-4ac}}{2a}=\frac{-b\pm\sqrt{\Delta}}{2a}\]

OpenStudy (anonymous):

oh you were dividing by 2 there which confused me a bit

OpenStudy (anonymous):

Oh i actually found out how to do this. thanks for the help

OpenStudy (kirbykirby):

:)